You are given the two vectors "\mathbf{a}=(1,2,1)" and "\mathbf{b}=(3,-4,2)" . Suppose that the pair of unit vectors perpendicular to both a and b are given by


"\pm \frac{1}{\sqrt{p}}(q,1,r),"

where p, q and r are some constants.



Determine the values of p, q and r.


1
Expert's answer
2021-10-04T16:18:22-0400

"c=\bold{a}\times\bold{b}=\begin{vmatrix}
 i & j & k \\
 1 & 2 & 1\\
 3 & -4 & 2
\end{vmatrix}\\
=i(2\cdot2-(-4)\cdot1)-j(1\cdot2-1\cdot3)+k(1\cdot(-4)-3\cdot2)=8i+j-10k\\
=\dfrac{1}{\sqrt{p}}(q,1,r)\\
8=\dfrac{q}{\sqrt{p}}\\
1=\dfrac{1}{\sqrt{p}}\\
\therefore
p=1\\
q=8\sqrt{1}=8\\
-10=\dfrac{r}{\sqrt{p}}\\
r=-10\sqrt{1}=-10"